HETAV PATEL

Grade 11,

A constant force F acts on a particle of mass 1kg moving with a velocity v, for one second. The distance moved in that time is . 0 F/2 v/2 v+F/2

A constant force F acts on a particle of mass 1kg moving with a velocity v, for one second. The distance moved in that time is .
  1. 0
  2. F/2
  3. v/2
  4. v+F/2

Grade:12th pass

1 Answers

Anisha
5 years ago
Mass=1kg
Acceleration=f/m=f/1=f
Initial velocity=v m/s
By newton's law :
S=ut+(1/2)at2 
Time=t=1sec
S=v+f/2 -:answer